OpenStudy (anonymous):

Is the line through the points R(−1, 3) and S(2, −7) parallel to the graph of the line given by the equation, 10x + 3y = 6? (1 point) No, the slopes have opposite signs. Yes, the slopes have the same sign and value. --->Yes, the lines both decrease to the right. No, the lines have unequal slopes.

OpenStudy (texaschic101):

10x + 3y = 6 3y = -10x + 6 y = -10/3x + 2 -- slope is -10/3 (-1,3) x1 = -1 and y1 = 3 (2,-7) x2 = 2 and y2 = -7 lets find the slope using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) m = (-7 - 3) / (2 -(- 1) m = -10/3 -- same slope yes, the slopes have the same sign and value.

OpenStudy (anonymous):

Thank you a lot that explanation helped :)

OpenStudy (texaschic101):

no problem :)

OpenStudy (amistre64):

another way to view this is; parallel lines have the same equations, but differ at most by a constant: 10x + 3y = 6 is parallel to 10x + 3y = k if we insert the values of R and S, and they equate, then they are on the same line that is parallel to the given line. does: 10(-1) + 3(3) = 10(2) + 3(-7) ??

OpenStudy (texaschic101):

I didn't think about it that way....thanks amistre, I learn something new everyday :)

OpenStudy (anonymous):

Yeah same here thanks